Find the value of sin2π/7+sin4π/7+sin8π/7.
Sir please help me out to do this question

Grade:12th pass

1 Answers

Sudha Haran
6 years ago
=sin2π/7+sin4π/7+sin8π/7=sin2π/7+sin3π7+sinπ/7=sinπ/7+sin2π/7+sin3π/7Now use formula of the continues series of Trigonometric equation=(sin3π/14*sin4π/14)/sinπ/14=(sinπ/2-sinπ/2)/2sinπ/14. =0